OpenStudy (anonymous):

Solve x3 = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

\[x ^{3}=\frac{ 1 }{ 8 }\] ?

OpenStudy (anonymous):

yes @KarlaKalurky

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

solve for the cube roots both sides

OpenStudy (mathmate):

@Marc1313 remember to use ^ to indicate exponential. Example, 3^x to mean \(3^x\)

OpenStudy (anonymous):

okY SO HOW DO I SOLVE IT

OpenStudy (anonymous):

sorry how do i solve it?

OpenStudy (anonymous):

i know the answer is 1/2

OpenStudy (anonymous):

but i dont know which one

OpenStudy (anonymous):

\[\sqrt[3]{x ^{3}}=\sqrt[3]{\frac{ 1 }{ 8 }}\] \[x^{3/3}=\frac{ 1 }{ 8 }^{1/3}\] \[x ^{1}=1/2\] \[x=1/2\]

OpenStudy (anonymous):

im not sure but i think we get \[\pm 1/2\] IF we get the SQUARE root of the value: 1/8 @mathmate is it 1/2 or \[\pm 1/2\]? :)

OpenStudy (anonymous):

i want to get the answer right and im not sure @KarlaKalurky

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@triciaal

OpenStudy (mathmate):

hints: \(\Large (\frac{1}{2})^3=?\) \(\Large (-\frac{1}{2})^3=?\)